Question #4 is the question I am requesting help on. It is exercise #4on chapter 16, in the book, “energy, environment, and climate” by Richard Wolfson.
media%2Fad0%2Fad0174c8-0b7e-420e-b58f-77
Answer #1

energy of a barrel of oil = 1700 kWh

20000 barrels / day = 20000/24 barrels/hour = 833.33 barrels/hour

Thus energy equivalent = 833.333 * 1700 = 1416.67 MW

at 40 % efficiency of electricity generation = 1416.67 MW * 0.4 = 566.67 MW excess energy per hour.